OpenStudy (anonymous):

In a relay race, the first runner of the team runs one-half of the distance plus one-half of a kilometre; and the second runner reached the finish line after running one-third of the remaining distance plus one-third of a kilometre. How long was the race?

OpenStudy (anonymous):

5 km check: (5/2+1/2) + (5/3+1/3) = 5

OpenStudy (anonymous):

let D = distance of the race. runner 1 ran D/2 + 1/2 runner 2 ran D/3 + 1/3 total distance = runner1 + runner2 D = (D/2+1/2) + (D/3+1/3) solving for D gives 5.